I am just learning about these math equations and I am stumped. Could you please help me? I'm not sure what formula I should use to figure this out.
if the perimeter of a rectangular pool is 60 meters and the length is 4 meters more than the width, how wide is the pool?
Thank you so much for your time.

Answer by checkley77(12844)

2L+2W=PERIMETER.
L=W+4
2(W+4)+2W=60
2W+8+2W=60
4W=60-8
4W=52
W=52/4
W=13 M. ANS FOR THE WIDTH.
L=13+4=17 M. ANS. FOR THE LENGTH.
PROOF:
2*17+2*13=60
34+26=60
60=60